Operations Report – Nov 15th

Crew 215 Operations Report 15-Nov-2019

SOL: 5

Name of person filing report: Shane Usher

Non-nominal systems:

Notes on non-nominal systems:

Generator: run

Hours run: 12 h 42 min

From what time last night: 1830

To what time this morning: 0712

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 68%

Diesel Reading – 49%

Station Propane Reading – 70%

Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 47

Water Meter (units): 1458114

Water (static tank) (gal): 392

Static to Loft Pump used – yes

Water in Green Hab (gal): 355.5

Water in Science Dome (gal): 0

Toilet tank emptied: yes

Deimos rover used: not at Hab

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Sojourner rover used: Assigned to Director

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Spirit rover used: yes

Hours: 111.0

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 67%

Currently charging: yes

Opportunity rover used: no

Hours: 66.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: yes

Hours: 117.1

Beginning charge: (Before EVA) : 100%

Ending charge: (On return from EVA, before charging): 80%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no

Reason for use: N/A

Oil Added? no

ATV Fuel Used (gal): 0

# Hours the ATVs were used today: 0

Notes on ATVs: Assigned to the director

HabCar used and why, where? With Director

CrewCar used and why, where? At Grand Junction awaiting Crew 216

General notes and comments: This report template is now produced through an Excel spreasheet template developed by Crew Engineer, Shane Usher.

Summary of internet: Functional

Summary of suits and radios: Functional

Summary of Hab operations: Kitchen benchtop installed. Sink plumbed by Shane Usher and Andrew Wheeler.

Summary of GreenHab operations: Atmospheric equalization to the green hab carried out during the afternoon, plants were watered (16L, 4.5 gal). The heater automatically came on at 5pm.

Summary of Science Dome operations: Washed and crushed gypsum samples, Re-bagged micrometeorite samples.

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of health and safety issues: Conducted a simulated medical emergency.

Questions, concerns and requests to Mission Support:

Copyright © The Mars Society. All rights reserved. | Main Site